# Disk Wind Feedback from High-mass Protostars. V. Application of Multi-Modal Machine Learning to Characterize Outflow Properties

### v1

@@ Abstract
Characterizing protostellar outflows is fundamental to understanding star formation feedback, yet traditional methods are often hindered by projection effects and complex morphologies. We present a multi-modal deep learning framework that jointly leverages spatial and spectral information from CO observations to infer outflow mass, inclination, and position angle ($PA$). Our model, trained on synthetic ALMA observations generated from 3D magnetohydrodynamic simulations, utilizes a cross-attention fusion mechanism to integrate morphological and kinematic features with probabilistic uncertainty estimation. Our results demonstrate that Vision Transformer architectures significantly outperform convolutional networks, showing remarkable robustness to reduced spatial resolution. Interpretability analysis reveals a physically consistent hierarchy: spatial features dominate across all parameters, whereas spectral profiles provide secondary constraints for mass and inclination. Applied to observational ALMA data, the framework delivers stable mass and $PA$ estimates with exceptionally tightly constrained inclination angles. This study establishes multi-modal deep learning as a powerful, interpretable tool for overcoming projection biases in high-mass star formation studies.
@@ I Introduction
Several theoretical models have been proposed to explain the launching of protostellar outflows, including the X-wind model (F. Shu et al., 1994) and the magnetic tower model (D. Lynden-Bell, 1996). Among these, the magnetocentrifugal disk wind model has emerged as the leading paradigm (R. D. Blandford & D. G. Payne, 1982; R. E. Pudritz & C. A. Norman, 1983). In this model, open magnetic field lines thread the surface layers of a Keplerian accretion disk, allowing gas to be centrifugally accelerated outward, forming a characteristic bipolar outflow composed of a fast, collimated jet and a slower, wide-angle wind. This mechanism has been extensively investigated through sophisticated magnetohydrodynamic (MHD) simulations incorporating diverse physical effects and environmental conditions (e.g., R. Ouyed et al., 2003; O. Gressel et al., 2015; G. Mattia & C. Fendt, 2020; A. L. Rosen, 2022).
A growing body of observational evidence, particularly from high-resolution facilities such as ALMA, supports this model in low- and intermediate-mass protostars. Observations reveal outflows with rotating molecular gas and conical wind morphologies, consistent with predictions from MHD disk wind theory (e.g., R. Launhardt et al., 2009; P. Bjerkeli et al., 2016; A. de Valon et al., 2022; R. Launhardt et al., 2023). These findings suggest that magnetocentrifugal launching may be a universal mechanism operating across a broad range of stellar masses.
@@ II.1 Magnetohydrodynamics Simulations
We utilize the 3D ideal MHD simulations from J. E. Staff et al. (2023), which model a disk wind outflow from a massive protostar. These simulations follow the Turbulent Core Accretion (TCA) model (C. F. McKee & J. C. Tan, 2003), with boundary conditions of the simulations tracking the self-consistent evolution of a protostar growing from 1 $M_{\odot}$ to over 24 $M_{\odot}$ within an initial 60 $M_{\odot}$ core embedded in a clump with $\Sigma_{\rm cl}=1\,\mathrm{g\,cm^{-2}}$ (Y. Zhang et al., 2014). The original simulation domain covers a single hemisphere of the outflow, extending from an injection surface 100 au above the disk to a distance of 26,500 au. For the analysis in this work, we mirror the simulation domain across the disk midplane ($x_{1}=0$) to construct a complete, bipolar outflow. A comprehensive description of the MHD simulation setup, code, and parameters can be found in J. E. Staff et al. (2023). In this study, we adopt the same simulation snapshots as those used by D. Xu et al. (2024), corresponding to protostellar masses of 2, 4, 8, 12, 16, 20, and 24 $M_{\odot}$.
@@ II.2 Synthetic Observations
Figure 1: Examples of synthetic outflows with different protostellar masses and inclination angles. Blue represents the blue-shifted outflow lobes ($50<v_{blue}<-3$ km s-1), while red represents the red-shifted lobes ($3<v_{red}<50$ km s-1).
@@
We use the publicly available radiative transfer code radmc-3d (C. P. Dullemond et al., 2012) to generate synthetic 12CO (2-1) spectra under non-LTE conditions. Level populations are computed based on local density and temperature, assuming that gas and dust temperatures are coupled. This assumption is valid for the dense regions ($n_{\rm H}\gtrsim 10^{5}\,{\rm cm}^{-3}$) that dominate our simulation domain, and dust temperatures are calculated as described in D. Xu et al. (2024). We also conduct radiative transfer from 20 different viewing angles, where the inclination angles are sampled by uniformly spacing the cosine of inclination angles between 0.025 and 0.975, ensuring even coverage over the spherical coordinate system.
@@
The input to the machine learning model includes a three-channel image: blue-shifted and red-shifted 12CO (2-1) integrated intensities, along with the total integrated intensity, each normalized between 0 and 1. The corresponding spectrum is clipped at $\pm$50 km/s and similarly normalized. To augment the dataset, we implement several strategies in the data loader: randomly shifting the outflow center in the image, swapping the blue- and red-shifted lobes in velocity, rotating the outflow by random position angles, adding Gaussian noise to the image, and injecting pattern noise from other outflows to mimic complex structures. The added pattern noise is at least a factor of three weaker than the main signal. In addition, we introduce random spectral shifts of $\pm$5 km/s to simulate velocity-centering uncertainties commonly encountered in real observations.
@@ II.3 Multi-Modal Machine Learning Framework
Besides the fiducial model training on all the physical parameters, we also conduct test model training which only includes certain physical setups. In particular, in these test models we exclude all samples with a protostellar mass of 12 $M_{\mathord{\odot}}$, and also exclude all samples with five specific inclination angles (with cosine values 0.125, 0.325, 0.525, 0.725 and 0.925), which means the training data does not cover any of these parameters, reducing the total training sample by 35%. This test model aims to show the how the model performs with limited dataset and, especially, the ability to predict properties of unseen data.
@@ III.1 Assessing Model Performance
Figure 4: Performance of the ViT_L_16 test model. See Appendix A (Figure 13) for a comparison of other model architectures. This model is trained on a restricted dataset that excludes samples with a protostellar mass of 12 $M_{\odot}$ and five specific inclination angles. Model performance is evaluated on outflows corresponding to the excluded mass and inclination angles. The top row compares the ground-truth values with the model predictions for protostellar mass, inclination angle, and position angle from left to right. The bottom row shows the probability distribution functions (PDFs) of the prediction errors, defined as the difference between the predicted and true values.
@@
We evaluate the performance of the trained multi-modal machine learning models using synthetic test datasets, beginning with an assessment of generalization under limited training coverage. In this setup, we exclude all samples with a protostellar mass of 12 $M_{\mathord{\odot}}$ and five specific inclination angles from the training phase, evaluating the models exclusively on these unseen parameter combinations. Figure 4 illustrates the performance of the ViT_L_16 architecture; for conciseness, the comparative results for all four model architectures are detailed in Figure 13 of Appendix A. A common limitation observed across all models is the discretization of predictions resulting from the sparse sampling of the physical parameter space. Consequently, predictions for the unseen 12 $M_{\mathord{\odot}}$ case tend to cluster around the nearest training grid points (e.g., 8, 16, or 24 $M_{\mathord{\odot}}$). However, the error distributions reveal distinct behaviors between architectures: ViT-based models systematically bias predictions toward 16 $M_{\mathord{\odot}}$, suggesting an ability to capture the continuous evolutionary progression of outflow morphology. In contrast, ResNet-based models exhibit broader, less structured biases, likely reflecting intrinsic morphological degeneracies exacerbated by beam convolution and noise. Despite these challenges in mass and inclination recovery, the position angle remains a robustly diagnosable parameter, with all four models achieving accurate recovery even for previously unseen outflow configurations.
@@
We next evaluate the fiducial models, which are trained on the full range of protostellar masses and inclination angles using the 1 kpc and 2 kpc datasets. Testing is conducted on the reserved 500 pc dataset, utilizing a fixed set of position angles sampled at 30 intervals to ensure consistent evaluation. Figure 5 summarizes the results for the ViT_L_16 model; for conciseness, comparative results for all four architectures are detailed in Figure 14 of Appendix A. Across the board, all models achieve accurate predictions for mass, inclination, and position angle, characterized by low bias and small uncertainties. Despite significant differences in model complexity—ranging from $\sim$27M parameters for ResNet50 to $\sim$307M for ViT_L_16-overall performance is broadly comparable, indicating that increasing model size does not yield uniform improvements. As shown in the Appendix, ResNet152 slightly underperforms the smaller ResNet50, suggesting potential over-parameterization relative to the training data. In contrast, the ViT-based models prove more robust to scaling, with ViT_L_16 outperforming the ResNet architectures, particularly in protostellar mass prediction.
@@
We then assess model robustness to resolution degradation by testing performance on synthetic data convolved with increasingly large Gaussian kernels. Figure 6 illustrates the performance of ViT_L_16, while comparative results for all four architectures are provided in Figure 15 of Appendix A. Although training augmentations included Gaussian smoothing with $\sigma$ between 0.75 and 2.0 (for $224\times 224$ inputs), we test well beyond this range. All models experience significant performance degradation when the kernel size exceeds $\sigma\sim 5$, reflecting the inevitable loss of morphological information at extreme smoothing levels. However, ViT_L_16 consistently provides the most stable position angle predictions under heavy blurring. Similarly, ViT_B_16 demonstrates superior generalization compared to ResNet-based models with equivalent parameter counts. These results suggest that transformer-based architectures are particularly well-suited for handling unseen observational conditions and degraded spatial information in protostellar outflow studies.
@@
Finally, we perform a controlled ablation test to isolate the contribution of the spectral modality. As detailed in § II.3, our training strategy involved randomly masking spectral segments, explicitly encouraging the network to prioritize spatial features. To evaluate this at inference, we disable the spectral branch entirely (setting inputs to zero) and compare performance against the standard multi-modal baseline. Figure 7 illustrates the results for ViT_L_16, while comparative plots for all four architectures are provided in Figure 16 of Appendix A. The results reveal architecture-dependent sensitivities. For ResNet-based models, including spectra improves inclination angle predictions but slightly degrades mass estimates, leaving position angles largely unaffected. Conversely, ViT_B_16 sees a modest improvement in mass prediction from spectral data with minimal impact on other parameters. Interestingly, for ViT_L_16, the inclusion of spectra results in a slight performance dip for both mass and inclination. This suggests that for this larger backbone, spectral features may introduce mild confusion under the current training setup, or potentially reflects test-set variance; regardless, position angle recovery remains robust across all configurations, confirming it is driven almost exclusively by spatial morphology. Overall, under moderate blurring conditions ($\sigma\lesssim 4$), model performance remains virtually unchanged whether spectral data is included or not. This behavior is consistent with our training strategy and corroborates the Integrated Gradients analysis in §III.2, confirming that spatial features are the dominant drivers of the predictions.

### v2

@@ Abstract
Characterizing protostellar outflows is fundamental to understanding star formation feedback, yet traditional methods are often hindered by projection effects and complex morphologies. We present a multi-modal deep learning framework that jointly leverages spatial and spectral information from CO observations to infer protostellar mass, inclination, and position angle ($PA$). Our model, trained on synthetic ALMA observations generated from 3D magnetohydrodynamic simulations, utilizes a cross-attention fusion mechanism to integrate morphological and kinematic features with probabilistic uncertainty estimation. Our results demonstrate that Vision Transformer architectures significantly outperform convolutional networks, showing remarkable robustness to reduced spatial resolution. Interpretability analysis reveals a physically consistent hierarchy: spatial features dominate across all parameters, whereas spectral profiles provide secondary constraints for mass and inclination. Applied to observational ALMA data, the framework delivers stable mass and $PA$ estimates with exceptionally tightly constrained inclination angles. This study establishes multi-modal deep learning as a powerful, interpretable tool for overcoming projection biases in high-mass star formation studies.
@@ I Introduction
Several theoretical models have been proposed to explain the launching of protostellar outflows, including the X-wind model (F. Shu et al., 1994) and the magnetic tower model (D. Lynden-Bell, 1996). Among these, the magnetocentrifugal disk wind model has emerged as the leading paradigm (R. D. Blandford & D. G. Payne, 1982; R. E. Pudritz & C. A. Norman, 1983). In this model, open magnetic field lines thread the surface layers of a Keplerian accretion disk, allowing gas to be centrifugally accelerated outward, forming a characteristic bipolar outflow composed of a fast, collimated jet and a slower, wide-angle wind. Once launched, this primary disk wind sweeps up and accelerates the surrounding ambient envelope material, producing the massive, entrained outflows that are most commonly observed in molecular tracers such as CO. In fact, it is only in the rare cases where rotating molecular gas is explicitly resolved that the primary disk wind itself is believed to be directly observed. This mechanism has been extensively investigated through sophisticated magnetohydrodynamic (MHD) simulations incorporating diverse physical effects and environmental conditions (e.g., R. Ouyed et al., 2003; O. Gressel et al., 2015; G. Mattia & C. Fendt, 2020; A. L. Rosen, 2022).
A growing body of observational evidence, particularly from high-resolution facilities such as ALMA, supports this model in low- and intermediate-mass protostars. Observations reveal outflows with rotating molecular gas and conical wind morphologies, notably in sources such as HH 212 (C.-F. Lee et al., 2021; J. A. López-Vázquez et al., 2024), L1527 (N. Sakai et al., 2017), TMC-1A (P. Bjerkeli et al., 2016), DG Tau B (A. de Valon et al., 2022), NGC 1333 IRAS 4C (Y. Zhang et al., 2018), and CB 26 (R. Launhardt et al., 2009, 2023). These findings, which are consistently matched with predictions from MHD disk wind theory, suggest that magnetocentrifugal launching may be a universal mechanism operating across a broad range of stellar masses. In the high-mass regime, a prominent example of such physics is the rotating bipolar outflow associated with Orion-KL Source I, where high-resolution observations have resolved a clear disk-wind structure (T. Hirota et al., 2017; J. A. López-Vázquez et al., 2020; M. Wright et al., 2024). This source provides critical evidence that magnetocentrifugal winds can successfully remove angular momentum and drive powerful outflows even in the presence of the intense radiative feedback characteristic of massive protostars.
@@ II.1 Magnetohydrodynamics Simulations
We utilize the 3D ideal MHD simulations from J. E. Staff et al. (2023), which model a disk wind outflow from a massive protostar. These simulations follow the Turbulent Core Accretion (TCA) model (C. F. McKee & J. C. Tan, 2003), with boundary conditions of the simulations tracking the self-consistent evolution of a protostar growing from 1 $M_{\odot}$ to over 24 $M_{\odot}$ within an initial 60 $M_{\odot}$ core embedded in a clump with $\Sigma_{\rm cl}=1\,\mathrm{g\,cm^{-2}}$ (Y. Zhang et al., 2014). This bounding pressure sets an initial core radius of $R_{c}=0.057$ pc. The core is threaded by a poloidal magnetic field with a total initial magnetic flux of $1~\mathrm{mG}\times R_{c}^{2}$. To manage the computational expense of resolving the inner disk over $\sim 10^{5}$ yr, the problem is divided into an evolutionary sequence of quasi-steady states. Because we directly analyze the simulation snapshots generated by J. E. Staff et al. (2023), our evolutionary sequence intrinsically follows the physical parameters summarized in their Table 1, which details the varying accretion rates, disk radii, and wind injection properties at each protostellar mass stage. The original simulation domain covers a single hemisphere of the outflow, extending from an injection surface located 100 au above the disk midplane to an outer boundary at a height of 26,500 au from the midplane. For the analysis in this work, we mirror the simulation domain across the disk midplane ($x_{1}=0$) to construct a complete, bipolar outflow. Consequently, the central $\pm 100$ au region is absent from the mirrored gas grid; however, because this missing region is entirely unresolved by the synthesized beam in our simulated ALMA observations, it does not impact the large-scale outflow morphology or the subsequent machine learning inference. A comprehensive description of the MHD simulation setup, code, and parameters can be found in J. E. Staff et al. (2023). In this study, we adopt the same simulation snapshots as those used by D. Xu et al. (2024), corresponding to protostellar masses of 2, 4, 8, 12, 16, 20, and 24 $M_{\odot}$.
@@ II.2 Synthetic Observations
Figure 1: Examples of synthetic 12CO (2-1) integrated intensity maps for outflows with different protostellar masses and inclination angles. Blue represents the blue-shifted outflow lobes ($50<v_{blue}<-3$ km s-1), while red represents the red-shifted lobes ($3<v_{red}<50$ km s-1).
@@
We use the publicly available radiative transfer code radmc-3d (C. P. Dullemond et al., 2012) to generate synthetic 12CO (2-1) spectra under non-LTE conditions. Level populations are computed based on local density and temperature, assuming that gas and dust temperatures are coupled. For these calculations, we assume a constant 12CO abundance of $10^{-4}$ relative to $H_{2}$ and a gas-to-dust mass ratio of 100. This assumption is valid for the dense regions ($n_{\rm H}\gtrsim 10^{5}\,{\rm cm}^{-3}$) that dominate our simulation domain, and dust temperatures are calculated as described in D. Xu et al. (2024). We also conduct radiative transfer from 20 different viewing angles, where the inclination angles are sampled by uniformly spacing the cosine of inclination angles between 0.025 and 0.975, ensuring even coverage over the spherical coordinate system. These values correspond to a range of inclination angles from approximately 12.8 to 88.6 relative to the line of sight.
@@
Similarly, Figure 2 displays the corresponding synthetic 12CO (2-1) spectra for these representative models. The spectral profiles reveal that the viewing geometry plays a dominant role in shaping the observable kinematics, particularly the high-velocity line wings. As shown in the right panels for a fixed mass of $12M_{\odot}$, varying the inclination angle ($\cos(i)$) significantly alters the spectral profile. Low-inclination (pole-on, $\cos(i)\approx 0.975$, $i\approx 12.8$) configurations yield significantly broader observable velocity wings because the outflowing gas is directed predominantly along the line of sight. In contrast, high-inclination (edge-on, $\cos(i)\approx 0.025$, $i\approx 88.6$) views result in much narrower velocity profiles, as the bulk of the outflow velocity lies in the plane of the sky. While the protostellar mass also influences the overall emission intensity and line shape, reflecting changes in the underlying wind power and envelope density as the system evolves (see left panels at $\cos(i)=0.425$, $i\approx 64.8$), its impact on the maximum extent of the line wings is secondary compared to the projection effects introduced by the inclination angle. These strong, inclination-dependent spectral variations provide the multi-modal network with critical kinematic constraints that complement the spatial morphologies in the integrated intensity maps.
The input to the machine learning model includes a three-channel image: blue-shifted and red-shifted 12CO (2-1) integrated intensities, along with the total integrated intensity, each normalized between 0 and 1. To account for missing flux near the systemic velocity, a common effect in ALMA observations, and to ensure the model is not overly sensitive to the precise definition of the central velocity, we employ a randomized velocity cutoff strategy. The blue- and red-shifted integrated intensity maps are generated using varying exclusion ranges ($|v-v_{\text{lsr}}|>3,5,8,$ or $10\text{ km/s}$), which mimics the filtering of low-velocity gas while preserving the large-scale structure of the outflow lobes. Visual inspection confirms that the resulting morphologies are robust across these ranges, allowing the network to learn consistent geometric cues despite variations in spectral coverage. The corresponding spectrum is clipped at $\pm$50 km/s and similarly normalized. To augment the dataset, we implement several strategies in the data loader: randomly shifting the outflow center in the image, swapping the blue- and red-shifted lobes in velocity, rotating the outflow by random position angles, adding Gaussian noise to the image, and injecting pattern noise from other outflows to mimic complex structures. The added pattern noise is at least a factor of three weaker than the main signal. In addition, we introduce random spectral shifts of $\pm$5 km/s to simulate velocity-centering uncertainties commonly encountered in real observations.
@@ II.3 Multi-Modal Machine Learning Framework
To ensure the network learns robust, physics-based representations rather than memorizing the discrete simulation grid, we implement a rigorous data splitting and dynamic augmentation strategy. To test scale invariance, our training and validation sets are strictly segregated by source distance: synthetic models placed at 1 kpc and 2 kpc are used exclusively for training, while models placed at 500 pc are reserved entirely for validation and testing. During training, we utilize an “on-the-fly” data loader that dynamically applies continuous augmentations to the input images and spectra at each epoch. These augmentations include adding random noise, convolving the images with random beam sizes ($\sigma$ between 0.75 and 2.0 pixels for $\times224$ inputs), applying random rotations ($0^{\circ}-360^{\circ}$), shifting the outflow central position to mimic real-world centering uncertainties, zooming in and out by up to 20%, and randomly swapping the blue- and red-shifted spatial lobes while simultaneously mirroring the corresponding 1D spectrum. This dynamic pipeline ensures the model—which contains $\sim$27 million parameters in the ResNet50 configuration and $\sim$307 million in the ViT_L_16 configuration—is exposed to a virtually infinite number of unique sample variations, heavily mitigating the risk of overfitting. Conversely, the validation and testing sets are evaluated strictly on raw, un-augmented data. The final testing set comprises 6,720 distinct samples, representing every combination of the 7 protostellar masses, 20 inclination angles, 4 velocity cutoff ranges ($\pm 3,\pm 5,\pm 8,\text{ and }\pm 10$ km/s), and 12 fixed position angles (sampled every $30^{\circ}$). The training and validation loss curves, which track the Negative Log-Likelihood (NLL) and Mean Squared Error (MSE), are presented in Appendix A. Our early-stopping protocol successfully captures the model weights at the global minimum before overfitting occurs; for instance, the fiducial ViT model achieved its best performance at epoch 392 with a validation NLL of -4.927 (comparable to its training NLL of -4.876) and a validation MSE of $7\times 10^{-6}$.
In addition to training our fiducial model on the complete physical parameter space, we construct a “restricted dataset” to conduct a controlled stress test of the network’s generalization capabilities. In this test setup, we deliberately withhold specific physical configurations from the training phase to evaluate the model’s performance on entirely unseen data. We specifically exclude all samples with a protostellar mass of $12M_{\odot}$, as this represents an intermediate evolutionary stage, providing an ideal baseline to test whether the network can physically interpolate between lower ($8M_{\odot}$) and higher ($16M_{\odot}$) masses. Similarly, we exclude five evenly spaced inclination angles ($82.8^{\circ},71.0^{\circ},58.3^{\circ},43.5^{\circ},$ and $22.3^{\circ}$, corresponding to cosine values of 0.125, 0.325, 0.525, 0.725, and 0.925) to evaluate the model’s geometric interpolation across the full viewing domain. By removing these parameters, we reduce the total training sample by 35%. This rigorous hold-out validation ensures that our assessment of the model’s predictive power (discussed in Section III.1) reflects true physical interpolation rather than overfitting to the discrete simulation grid.
@@ III.1 Assessing Model Performance
Figure 4: Performance of the ViT_L_16 test model. See Appendix B (Figure 16) for a comparison of other model architectures. This model is trained on a restricted dataset that excludes samples with a protostellar mass of 12 $M_{\odot}$ and five specific inclination angles. Model performance is evaluated on outflows corresponding to the excluded mass and inclination angles. The top row compares the ground-truth values with the model predictions for protostellar mass, inclination angle, and position angle from left to right. The bottom row shows the probability distribution functions (PDFs) of the prediction errors, defined as the difference between the predicted and true values.
@@
We evaluate the performance of the trained multi-modal machine learning models using synthetic test datasets, beginning with an assessment of generalization under limited training coverage. In this setup, we exclude all samples with a protostellar mass of 12 $M_{\mathord{\odot}}$ and five specific inclination angles from the training phase, evaluating the models exclusively on these unseen parameter combinations. This restricted setup serves as a rigorous test for generalization; by evaluating the model on physical configurations that were entirely excluded from the training set across all simulated distances, we can distinguish between the network’s ability to interpolate physical trends and the mere memorization of pixel patterns from high-resolution counterparts. Figure 4 illustrates the performance of the ViT_L_16 architecture; for conciseness, the comparative results for all four model architectures are detailed in Figure 16 of Appendix B. A common limitation observed across all models is the discretization of predictions resulting from the sparse sampling of the physical parameter space. Consequently, predictions for the unseen 12 $M_{\mathord{\odot}}$ case tend to cluster around the nearest training grid points (e.g., 8, 16, or 24 $M_{\mathord{\odot}}$). However, the error distributions reveal distinct behaviors between architectures: ViT-based models systematically bias predictions toward 16 $M_{\mathord{\odot}}$, suggesting an ability to capture the continuous evolutionary progression of outflow morphology. In contrast, ResNet-based models exhibit broader, less structured biases, likely reflecting intrinsic morphological degeneracies exacerbated by beam convolution and noise. We note that these structured biases are specifically a consequence of the sparse parameter coverage in this restricted test setup; as demonstrated in the following subsection, these degeneracies are effectively resolved when the architectures are trained on the full physical parameter space. Despite these challenges in mass and inclination recovery, the position angle remains a robustly diagnosable parameter, with all four models achieving accurate recovery even for previously unseen outflow configurations.
@@
We next evaluate the fiducial models, which are trained on the full range of protostellar masses and inclination angles using the 1 kpc and 2 kpc datasets. Testing is conducted on the reserved 500 pc dataset, utilizing a fixed set of position angles sampled at 30 intervals to ensure consistent evaluation. Figure 5 summarizes the results for the ViT_L_16 model; for conciseness, comparative results for all four architectures are detailed in Figure 17 of Appendix B. Across the board, all models achieve accurate predictions for mass, inclination, and position angle, characterized by low bias and small uncertainties. Despite significant differences in model complexity, ranging from $\sim$27 million parameters for ResNet50 to $\sim$307 million for ViT_L_16, overall performance is broadly comparable, indicating that increasing model size does not yield uniform improvements. As shown in the Appendix, ResNet152 slightly underperforms the smaller ResNet50, suggesting potential over-parameterization relative to the training data. In contrast, the ViT-based models prove more robust to scaling, with ViT_L_16 outperforming the ResNet architectures, particularly in protostellar mass prediction.
@@
We then assess model robustness to resolution degradation by testing performance on synthetic data convolved with increasingly large Gaussian kernels. Figure 6 illustrates the performance of ViT_L_16, while comparative results for all four architectures are provided in Figure 18 of Appendix B. Although training augmentations included Gaussian smoothing with $\sigma$ between 0.75 and 2.0 (for $224\times 224$ inputs), we test well beyond this range. All models experience significant performance degradation when the kernel size exceeds $\sigma\sim 5$, reflecting the inevitable loss of morphological information at extreme smoothing levels. However, ViT_L_16 consistently provides the most stable position angle predictions under heavy blurring. Similarly, ViT_B_16 demonstrates superior generalization compared to ResNet-based models with equivalent parameter counts. These results suggest that transformer-based architectures are particularly well-suited for handling unseen observational conditions and degraded spatial information in protostellar outflow studies.
@@
Finally, we perform a controlled ablation test to isolate the contribution of the spectral modality. As detailed in § II.3, our training strategy involved randomly masking spectral segments, explicitly encouraging the network to prioritize spatial features. To evaluate this at inference, we disable the spectral branch entirely (setting inputs to zero) and compare performance against the standard multi-modal baseline. Figure 7 illustrates the results for ViT_L_16, while comparative plots for all four architectures are provided in Figure 19 of Appendix B. The results reveal architecture-dependent sensitivities. For ResNet-based models, including spectra improves inclination angle predictions but slightly degrades mass estimates, leaving position angles largely unaffected. Conversely, ViT_B_16 sees a modest improvement in mass prediction from spectral data with minimal impact on other parameters. Interestingly, for ViT_L_16, the inclusion of spectra results in a slight performance dip for both mass and inclination. This suggests that for this larger backbone, spectral features may introduce mild confusion under the current training setup, or potentially reflects test-set variance; regardless, position angle recovery remains robust across all configurations, confirming it is driven almost exclusively by spatial morphology. Overall, under moderate blurring conditions ($\sigma\lesssim 4$), model performance remains virtually unchanged whether spectral data is included or not. This behavior is consistent with our training strategy and corroborates the Integrated Gradients analysis in §III.2, confirming that spatial features are the dominant drivers of the predictions.